# Strategic Communication and Language Bias in Multi-Agent LLM Coordination

### v1

@@ 4 Conclusion and Future Directions
This study presents a systematic investigation into how communication influences the strategic behavior of LLM-based agents in multi-agent environments. By extending the FAIRGAME framework with a communication layer, we analyzed the effects of message exchange in two canonical game-theoretic scenarios, the Prisoner’s Dilemma and the Battle of the Sexes, across multiple languages, personality profiles, and game configurations. Comparing Llama 4 Maverick and GPT-4o, we found that communication substantially alters agent behavior: in the Prisoner’s Dilemma, a social dilemma involving trust and self-interest, communication fosters cooperation and improves outcomes; by contrast, in the Battle of the Sexes, a coordination game, communication unexpectedly reduces the level of alignment between agents.
Our analysis reveals that communicative behavior is deeply context-sensitive. Message length, lexical choices, and strategic alignment vary with game type, personality dynamics, and language. These patterns underscore the dual role of language in LLM interactions: as a vehicle for coordination and as a potential vector for emergent biases.
This work lays the groundwork for a broader research agenda at the intersection of language, strategy, and AI coordination. Future investigations should extend to a wider set of languages, especially low-resource or typologically diverse ones, to assess the generalizability and fairness of communicative strategies. Clusters of related languages may illuminate whether linguistic similarity drives behavioral convergence, or whether model-specific traits dominate.
Deeper semantic analysis of agent communication remains a critical next step. While message length and frequency offer useful signals, uncovering strategic intent, negotiation dynamics, and emergent alignment will require more sophisticated tools. Advances in LLM-based evaluators and conversation modeling could enable scalable, nuanced interpretation of agent interactions—though such methods must also confront issues like evaluator bias.
Ultimately, understanding how LLMs reason, negotiate, and cooperate through language in complex social settings is key to building robust and trustworthy multi-agent systems. This study marks a step in that direction, emphasizing that communication is not an auxiliary feature but a central mechanism shaping AI behavior.
These findings inform the design of safer and more interpretable multi-agent AI systems, especially in applications where strategic alignment and fairness are critical [15, 19, 21].

### v2

@@ 4 Conclusion and Future Directions
This study systematically investigates how communication influences the strategic behavior of LLM-based agents in multi-agent environments. By extending the FAIRGAME framework with a communication layer, we analyzed the effects of message exchange in two canonical game-theoretic scenarios, the Prisoner’s Dilemma and the Battle of Sexes, across multiple languages, personality profiles, and game configurations. Comparing Llama 4 Maverick and GPT-4o, we found that communication substantially alters agent behavior: in the Prisoner’s Dilemma, a social dilemma involving trust and self-interest, communication fosters cooperation and improves outcomes; by contrast, in the Battle of Sexes, a coordination game, communication unexpectedly reduces the level of alignment between agents.
Our analysis reveals that communicative behavior is deeply context-sensitive. Message length, lexical choices, and strategic alignment vary with game type, personality dynamics, and language. These patterns underscore the dual role of language in LLM interactions: as a vehicle for coordination and as a potential vector for emergent biases. This work lays the groundwork for a broader research agenda at the intersection of language, strategy, and AI coordination. Future investigations should extend to a wider set of languages, especially low-resource or typologically diverse ones, to assess the generalizability and fairness of communicative strategies. Clusters of related languages may illuminate whether linguistic similarity drives behavioral convergence, or whether model-specific traits dominate. Deeper semantic analysis of agent communication remains a critical next step. While message length and frequency offer useful signals, uncovering strategic intent and negotiation dynamics will require more sophisticated tools. Advances in LLM-based evaluators and conversation modeling could enable scalable, nuanced interpretation of agent interactions, though such methods must also confront issues like evaluator bias. Ultimately, understanding how LLMs reason, negotiate, and cooperate through language in complex social settings is key to building robust and trustworthy multi-agent systems. This study marks a step in that direction, emphasizing that communication is not an auxiliary feature but a central mechanism shaping AI behavior. These findings inform the design of safer and more interpretable multi-agent AI systems, especially in applications where strategic alignment and fairness are critical [13, 17, 19].